find three consecutive positive integers, such that the square of the smallest exceeds the largest by 10
Answer by mananth(16946)

let the integers be x,x+1 & x+2
x^2= (x+2)+10
x^2=x+12
x^2-x-12=0
x^2-4x+3x-12=0
x(x-4)+3(x-4)=0
(x-4)(x+3)=0
x=4 OR -3
positive integer = 4
4,5,6